7. The value of a lathe originally valued at K3000 depreciates 15%per annum. Calculate its value after 4 years. The machine is sold when its value is less than K550. After how many years is the lathe sold? (hint : use GP)
8. On commencing employment a man is paid a salary of £7200 per annum and receives
annual increments of £350. Determine his salary in the 9th year and calculate the total
he will have received in the first 12 years.
9. A hire tool firm finds that their net return from hiring tools is decreasing by 10% per annum. If their net gain on a certain tool this year is £400, find the possible total of all future profits
from this tool (assuming the tool lasts for ever).
10. K1200 is invested at 12.5% simple interest. How much will have accrued after 3 years?
11. K10000 was compounded at 12.5% and amounted to K52015.80. How many years did it take?

Answer by Boreal(15235)

7. 3000*(0.85)^4*3000, the 85% is 100%-15% depreciation. This is K1566.02
550=3000(.85)^t
0.18333=.85^t
ln both sides
-1.70=t ln .85, round at end
divide both sides by ln .85
t=10.44 years
8. in the 9th year, he has received 8 increments of 350 pounds or 2800.
his salary is 10,000 pounds after 9 years and 11,150 after 12 years
Sum12 is (n/2)(2 a1+(n-1)d)=6(14400+11(350))=109,500 pounds
9.Geometric series r=0.9, sum is a/1-r=400/.1=4000 pounds
10. I=prt=1200*1/8*3=$450 simple interest accruing after 3 years.
11. assuming compounding is annual, 52015.80=10000(1+.125)^n
5.20128=1.125^n
ln both sides
then divide by ln 1.125
1.6489=t 1n 1.125 round at end
14.00 years if compounded annually.